A television network has a choice of 11 new shows for four consecutive timeslots. How many ways can four shows be chosen without considering the age of the viewers or the popularity of the timeslots
Answer by Boreal(15235)

11C4=11!/4!7!
=11*10*9*8*7!/7!*4!
=110*72/24
=110*3
=330 ways